# The Collection of a Human Robot Collaboration Dataset for Cooperative Assembly in Glovebox Environments

### v1

@@ Abstract
Industry 4.0 introduced AI as a transformative solution for modernizing manufacturing processes. Its successor, Industry 5.0, envisions humans as collaborators and experts guiding these AI-driven manufacturing solutions. Developing these techniques necessitates algorithms capable of safe, real-time identification of human positions in a scene, particularly their hands, during collaborative assembly. Although substantial efforts have curated datasets for hand segmentation, most focus on residential or commercial domains. Existing datasets targeting industrial settings predominantly rely on synthetic data, which we demonstrate does not effectively transfer to real-world operations. Moreover, these datasets lack uncertainty estimations critical for safe collaboration. Addressing these gaps, we present HAGS: Hand and Glove Segmentation Dataset. This dataset provides 1200 challenging examples to build applications toward hand and glove segmentation in industrial human-robot collaboration scenarios as well as assess out-of-distribution images, constructed via green screen augmentations, to determine ML-classifier robustness. We study state-of-the-art, real-time segmentation models to evaluate existing methods. Our dataset and baselines are publicly available [^1] https://dataverse.tdl.org/dataset.xhtml?persistentId=doi:10.18738/T8/85R7KQ  [^2] https://github.com/UTNuclearRoboticsPublic/assembly_glovebox_dataset.
@@ 1 Introduction
Gloveboxes are self-contained spaces that allow workers to handle hazardous materials through gloves affixed to sealed portholes (see Figure 1.) This setup protects operators from exposure and prevents unfiltered material releases into the environment (13). Workers and researchers that work with hazardous materials in gloveboxes face issues such as ergonomic injuries and potential hazardous exposure via glove tear. The use of robots in these environments can mitigate many of these problems.
@@
Figure 1: Left: (a) Example of a glovebox suite from (14). Right: (b,c) Examples of RGB images from the HAGs dataset (31) with corresponding pixel-wise labels.
@@
Recent successes in the machine learning community have inspired robotics researchers to develop large-scale datasets aimed at achieving breakthroughs similar to ImageNet (12) for robotic research (5; 9). While there are innumerable possibilities for robotics applications, previous datasets prioritize instruction following in residential environments (5; 9). Unfortunately, these datasets often overlook elements of Human-Robot Interaction (HRI), including Human-Robot Collaboration (HRC): where human and robotic agents work together toward a shared goal. In particular, this behavior is desired in manufacturing tasks for collaborative assembly.
@@
To perform collaborative assembly tasks safely, robots must understand where human operators’ hands are within a shared task space. This requires active safety systems which rely on hand segmentation algorithms to avoid or interact with human collaborators. Despite the plethora of openly available hand segmentation datasets, most do not prioritize operating in hazardous or industrial settings. Rather, these datasets leverage web-sourced data which are biased toward easily accessible objects and environments, e.g., common household items in residential settings (9). The public datasets that do target industrial domains often suffer from being either: (1) small-scale and lacking human subject diversity (30; 32) or (2) generated as synthetic data (18).
@@
These shortcomings are highly consequential for active safety systems. For instance, most real-time segmentation algorithms leverage convolutional neural networks (CNNs) for hand classification. However, these architectures may over-rely on pixel color values for classification. Thus underrepresentation in these datasets may unnecessarily imperil people of color.
@@
In an effort to motivate work toward this overlooked task space, we present HAGS: the Hand and Glove Segmentation dataset (IRB ID: STUDY00003948) which is publicly available  [^3] https://dataverse.tdl.org/dataset.xhtml?persistentId=doi:10.18738/T8/85R7KQ. This dataset contributes the following:
@@
    The inclusion of twelve diverse participants to maximize the relevance of the ungloved data.
@@ 2 Related Work
There are numerous hand segmentation datasets; however, most focus on aspects of daily life with activities such as cooking or playing cards (1; 21; 23). For a fuller accounting of these priors work see Appendix A.2. Although expansive, these datasets are not adequate for situations such as collaborative tasks with robots. Thus, we focus our dataset comparisons on the most relevant works in industrial domains (see Table 1.)
@@
WorkingHands, 2019 (32) CD Assembly 7.8k - Pixel 14
MECCANO, 2020 (28) C Assembly - 20 Bound. Box 21
HRC, 2022 (30) C Assembly 1.3k 2 Pixel 5
HaDR, 2023 (18) CD Manipulation 117k - Pixel -
HAGS (ours), 2024 (31) C Assembly 1.7k 10 Pixel 2
@@
Task Comparisons. Most industrial datasets are not tailored for HRC barring recent work on uncertainty estimation for segmentation tasks pertaining to collaborative assembly (30). Unfortunately, this dataset is derived from only two participants. In an effort to expand on this work, we collect additional data on twelve participants for two new assembly tasks while incorporating glovebox environment data. Supplementing glovebox data is crucial because this metallic environment presents unique challenges to segmentation such as shine and reflection.
@@
The HaDR dataset (18) features robotic arms and closely resembles our goal to develop robust hand segmentation algorithms. However, no specific HRC task is defined in their dataset and it is entirely synthetic. Similarly, many industrial datasets synthetically augment background color and texture to assess robustness in segmentation models (18; 32). Additionally, while these datasets contain a plethora of tool and item classes, they do not distinguish right- and left-hand information in their labeling, which presents a challenge to extending to future HRI research and applications.
@@
Participant Diversity. Well-constructed datasets are crucial to machine learning models’ reliability and task performance. In the context of active safety systems, participant diversity is of utmost importance. Our work surpasses previous industrial datasets (2; 6; 15; 25; 30) in this regard by including twelve diverse participants. Please refer to Appendix A.1 for additional details.
@@
Color Invariance. Several datasets leverage depth data to reduce the influence of lighting and skin coloration on their segmentation models (18; 32). However, multi-modal models may still over-rely on RGB features. Other work seeks to completely remove this bias by creating color-agnostic datasets but potentially risks generalizable segmentation performance by forgoing rich RGB signals (20).
@@
Alternative methods used to address color-invariant segmentation include augmenting existing works with simulated data (32) or generating fully synthetic training datasets (18). Although most synthetic datasets strive to capture features of real-world data, others aim to build texture and lighting invariance into their models via unrealistic data augmentations (18).
@@
Contributions of HAGS. Our work addresses the limitations of existing industrial domain datasets by focusing on a previously underserved area: glovebox environments. We present a significantly large and diverse dataset, comprising real images from ten participants, in contrast to the synthetic images commonly used in other studies. Recognizing the potential application of this technology in active safety systems for human-robot interactions, we have meticulously developed challenging examples using out-of-distribution (OOD) scenarios, such as green screens with distracting images and hands instead of gloves. Furthermore, we encourage the evaluation of uncertainty quantification (UQ) metrics, such as expected calibration error, to enhance safety information. Our findings indicate that prior works do not transfer effectively to our challenging dataset, underscoring the need for further efforts targeting industrial domains.
@@ 3 HAGS Dataset
This section provides an overview of the dataset collection, preparation, and annotation processes for HAGS. Appendix A.1 documents the complete Datasheet (17).
@@ 3.1 Data Collection
Videos are collected in a standard Department of Energy (DOE) glovebox. Two camera angles are provided per video: one 1080p GoPro from a bird’s-eye view and one 1080p Intel RealSense Development Kit Camera recording the right side of the participant. Twelve participants are included in the study with 16 videos each totaling more than 9 hours of content. Normally distributed frames are collected and annotated from each video, amounting to over 1440 frames. The Unreal Robotics UR3e robot arm, with an attached gripper for handling objects, is used to aid the human subject. The robot is pre-programmed to assist the human participant with sequential, assembly tasks.
@@ 3.2 Surrogate Joint-Assembly Tasks
Green screen. A green screen was placed on the bottom of the glovebox for half of the participant videos. This green screen was used to later apply synthetic textures and colors to the background of the real-world images (see Fig. 2). The script incorporated in our code repository for generating synthetic images leverages Google Image searches. Notably, these images are not stored within our dataset; rather, we provide the capability for other researchers to independently construct analogous datasets. Frames extracted from videos containing a green screen background are placed into an OOD test set.
Gloves. Despite operating in a glovebox, half of the recorded videos feature ungloved hands (see the right side of Fig. 1.) We incorporated this additional OOD scenario due to the risk of glove tears. Thus our dataset accounts for rare scenarios where the active safety system must still ensure safe HRC despite skin exposure. These videos are also placed in an OOD test set.
@@ 3.5 Data Annotation
Three classes were assigned to each image: left-hand, right-hand, and background. Human annotators were instructed to annotate each hand from the tip to the wrist, and provide their best estimate of the wrist location when the subject was wearing gloves. Using LabelStudio (34), annotators provided key point prompts to MobileSAM (40), a segmentation model which provided a coarse label for annotators to refine. Four of the researchers performed annotations. Two annotators labeled each image to develop inter-annotator agreement (IAA) for label quality. We calculate IAA in two ways: (1) as the average Cohen’s Kappa (0.916) and (2) as the average IOU (0.957) between annotator-provided labels across the full dataset , indicating strong agreement. Each frame’s annotations were converted to a single PNG file, where the three classes were recorded: left-hand, right-hand, background.
@@ 4 Experiments
We conducted two experiments to assess the challenging nature of our dataset. Experiment A is a transfer learning experiment designed to demonstrate the limitations of prior works, listed in Table 1, when applied to our task space. This experiment highlights the lack of sufficient transferability of pretraining on existing datasets to our specific domain. Experiment B studies uncertainty quantification (UQ) involving ID and OOD testing using the HAGS dataset. The OOD frames for all participants include scenarios where gloves are not worn or a green screen is placed in the background. For the green screen frames, we overlay images of different objects to create diverse and challenging testing conditions. This experiment assesses the robustness and reliability of our models in handling diverse, challenging scenarios.
@@ 4.1 Experimental Setup
For training all models in both experiments, we input 256x256 image sizes and used the Adam optimizer with a learning rate of 1e-3 for the UNet and BiSeNetv2 architectures, and 8e-4 for MobileSAM training. To train or perform inference with MobileSAM, a visual prompt is first required. We adopt a simple approach of selecting a bounding box prompt that contained the whole image. Dropout of p=0.1 was utilized for training of all architectures. We use a variety of data augmentations applied to our training set, including: resize, color jitter, advanced blur, Gaussian noise, random rotate 90 (p=0.5).
Experiment A. To train, we use an internal cluster equipped with eight NVIDIA RTX A5000 GPUs. The models undergo initial pretraining using the WorkingHands (32), HaDr (18), and HRC (30) datasets. Following this pretraining phase, we fine-tune the models on a subset (10%) of our ID HAGS dataset. This approach is designed to test the assertion of prior work that strong performance can be achieved with minimal real-world examples when leveraging their synthetic datasets. We perform this fine tuning for 20 epochs with batch sizes of 64 and 34 for UNet and BiSeNet respectively. We did not include the MECCANO (28) dataset in our study because it only provides bounding box labels, and our task requires pixel-level segmentation. We evaluate the models on both ID and OOD data from the HAGS dataset.

### v2

@@ Abstract
Industry 4.0 introduced AI as a transformative solution for modernizing manufacturing processes. Its successor, Industry 5.0, envisions humans as collaborators and experts guiding these AI-driven manufacturing solutions. Developing these techniques necessitates algorithms capable of safe, real-time identification of human positions in a scene, particularly their hands, during collaborative assembly. Although substantial efforts have curated datasets for hand segmentation, most focus on residential or commercial domains. Existing datasets targeting industrial settings predominantly rely on synthetic data, which we demonstrate does not effectively transfer to real-world operations. Moreover, these datasets lack uncertainty estimations critical for safe collaboration. Addressing these gaps, we present HAGS: Hand and Glove Segmentation Dataset. This dataset provides challenging examples to build applications toward hand and glove segmentation in industrial human-robot collaboration scenarios as well as assess out-of-distribution images, constructed via green screen augmentations, to determine ML-classifier robustness. We study state-of-the-art, real-time segmentation models to evaluate existing methods. Our dataset and baselines are publicly available.
keywords:
Semantic segmentation, Machine Learning, Glovebox, Dataset, Out-of-Distribution Classification
@@ 1 Introduction
Gloveboxes are self-contained spaces that allow workers to handle hazardous materials using gloves affixed to sealed portholes (see Figure 1.) This setup protects operators from exposure and prevents unfiltered material releases into the environment (DOE, 2003). Workers and researchers that handle hazardous materials in gloveboxes face issues such as ergonomic injuries and potential hazardous exposure via glove tear. The use of robots in these environments can mitigate many of these problems. This paper seeks to improve glovebox manufacturing techniques for applications involving Special Nuclear Materials (SNMs) (Bean et al., 2010) to minimize radiation exposure to personnel. SNM handing is important; for example the production of Mo-99 decay products, used in the majority of medical diagnostic procedures, involves the use of highly enriched uranium (Council, 2009). SNM manufacturing applications include medical isotope production (Council, 2009), DoD applications (Pryor and Williams, 2013), radiological imaging (Hashem et al., 2018), disposition of spent nuclear fuel (Ray et al., 2015), and handling irradiated samples for neutron activation analysis (Hashem et al., 2017). Current manufacturing procedures potentially expose personnel to extremely high radiation doses, and robotics has the potential to reduce the time and increase the distance between radioactive sources and personnel. Ideally applications are fully automated, but niche applications and experimental activities related to handling SNMs will likely require humans to remain - at least partially - in the loop. The safety standards a robotic system must meet in all of these applications is exceptionally stringent.
@@
Figure 1: Left: (a) Example of a glovebox suite from (DOE, 2020). Right: (b,c) Examples of RGB images from the HAGs dataset (Sharma et al., 2024) with corresponding pixel-wise labels.
@@
Recent successes in the machine learning community have inspired robotics researchers to develop large-scale datasets aimed at achieving breakthroughs similar to ImageNet (Deng et al., 2009) for robotic research (Brohan et al., 2023; Collaboration, 2023). While there are innumerable possibilities for robotics applications, previous datasets prioritize instruction following in residential environments (Brohan et al., 2023; Collaboration, 2023). Unfortunately, these datasets often overlook elements of Human-Robot Interaction (HRI), including Human-Robot Collaboration (HRC): where human and robotic agents work together toward a shared goal. In particular, this behavior is desired in manufacturing tasks for collaborative assembly.
@@
To perform collaborative assembly tasks safely, robots must understand where human operators’ hands are within a shared task space. This requires active safety systems which rely on hand segmentation algorithms to avoid or interact with human collaborators. Despite the plethora of openly available hand segmentation datasets, most do not prioritize operating in hazardous or industrial settings. Rather, these datasets leverage web-sourced data that are biased toward easily accessible objects and environments, e.g., common household items in residential settings (Collaboration, 2023). The public datasets that do target industrial domains often suffer from being either: (1) small-scale and lacking human subject diversity (Sajedi et al., 2022; Shilkrot et al., 2019) or (2) generated as synthetic data (Grushko et al., 2023).
@@
These shortcomings are highly consequential for active safety systems. For instance, most real-time segmentation algorithms leverage convolutional neural networks (CNNs) for hand classification. However, these architectures may over-rely on pixel color values for classification (Singh et al., 2020). Thus, underrepresentation in these datasets may unnecessarily imperil people of color.
@@
In an effort to motivate work toward this overlooked task space, we present HAGS: the Hand and Glove Segmentation dataset (IRB ID: STUDY00003948) which is publicly available within our dataset repository. This dataset contributes the following:
@@
    The inclusion of ten diverse participants to maximize the relevance of ungloved data.
@@ 2 Related Work
There are numerous hand segmentation datasets; however, most focus on aspects of daily life with activities such as cooking or playing cards (Bambach et al., 2015; Khan and Borji, 2018; Li et al., 2020). For a more comprehensive account of these prior works, refer to our Overview of Hand Segmentation Datasets in the Supplemental Material. Although expansive, these datasets are not adequate for situations such as collaborative tasks with robots. Thus, we focus our dataset comparisons on the most relevant works in industrial domains (see Table 1.)
@@
WorkingHands, 2019 (Shilkrot et al., 2019) CD Assembly 7.8k - Pixel 14
MECCANO, 2020 (Ragusa et al., 2020) C Assembly - 20 Bound. Box 21
HRC, 2022 (Sajedi et al., 2022) C Assembly 1.3k 2 Pixel 5
HaDR, 2023 (Grushko et al., 2023) CD Manipulation 117k - Pixel -
HAGS (ours), 2024 (Sharma et al., 2024) C Assembly 1.7k 10 Pixel 2
@@
Task Comparisons. Most industrial datasets are not tailored for HRC barring recent work on uncertainty estimation for segmentation tasks pertaining to collaborative assembly (Sajedi et al., 2022). Unfortunately, this dataset is derived from only two participants. In an effort to expand on this work, we collect additional data on ten participants for two new assembly tasks while incorporating glovebox environment data. Supplementing glovebox data is crucial because this metallic environment presents unique challenges to segmentation such as shine and reflection.
@@
The HaDR dataset (Grushko et al., 2023) features robotic arms and closely resembles our goal to develop robust hand segmentation algorithms. However, no specific HRC task is defined in their dataset and it is entirely synthetic. Similarly, many industrial datasets synthetically augment background color and texture to assess robustness in segmentation models (Grushko et al., 2023; Shilkrot et al., 2019). Additionally, while these datasets contain a plethora of tool and item classes, they do not distinguish right and left-hand information in their labeling, which presents a challenge to extending to future HRI research and applications.
@@
Participant Diversity. Well-constructed datasets are crucial to machine learning models’ reliability and task performance. In the context of active safety systems, participant diversity in machine learning datasets is of utmost importance. Our work supplements previous industrial datasets (Baraldi et al., 2014; Cai et al., 2017; Fathi et al., 2011; Likitlersuang and Zariffa, 2018; Sajedi et al., 2022) in this regard by including ten diverse participants. Please refer to our datasheet provided in the Supplemental Materials for additional details.
@@
Color Invariance. Several datasets leverage depth data to reduce the influence of lighting and skin coloration on their segmentation models (Grushko et al., 2023; Shilkrot et al., 2019). However, multi-modal models may still over-rely on RGB features (Singh et al., 2020). Other work seeks to completely remove this bias by creating color-agnostic datasets but potentially risks generalizable segmentation performance by forgoing rich RGB signals (Kang et al., 2018).
@@
Alternative methods used to address color-invariant segmentation include augmenting existing work with simulated data (Shilkrot et al., 2019) or generating fully synthetic training datasets (Grushko et al., 2023). Although most synthetic datasets strive to capture features of real-world data, others aim to build texture and lighting invariance into their models via unrealistic data augmentations (Grushko et al., 2023).
@@
Contributions of HAGS. Our work addresses the limitations of existing industrial domain datasets by focusing on a previously underserved area: glovebox environments. We present a sizable and diverse dataset comprising real images of ten participants, in contrast to the synthetic images commonly used in other studies. Recognizing the potential application of this technology in active safety systems for human-robot interactions, we have meticulously developed challenging examples using out-of-distribution (OOD) scenarios, such as green screens with distracting images and hands instead of gloves. Furthermore, we encourage the evaluation of uncertainty quantification (UQ) metrics, such as expected calibration error, to enhance safety information. Our findings indicate that prior works do not transfer effectively to our challenging dataset, underscoring the need for further efforts targeting industrial domains.
@@ 3 HAGS Dataset
This section provides an overview of the dataset collection, preparation, and annotation processes for HAGS. The Datasheet (Gebru et al., 2021) provided in the Supplemental Materials provides a more thorough accounting of the dataset.
@@ 3.1 Data Collection
Videos are collected in a standard Department of Energy (DOE) glovebox. Two camera angles are provided per video: one 1080p GoPro from a bird’s-eye view and one 1080p Intel RealSense Development Kit Camera recording the right side of the participant. Ten participants are included in the study with 16 videos each totaling more than 9 hours of content. Normally distributed frames are collected and annotated from each video, amounting to over 1440 frames. The Unreal Robotics UR3e robot arm, with an attached gripper for handling objects, is used to aid the human subject. The robot is pre-programmed to assist the human participant with sequential, assembly tasks.
@@ 3.2 Surrogate Joint-Assembly Tasks
The selection of Jenga as a surrogate task supports two critical areas of interest. First, Jenga presents a challenging and generalized pick-and-place assembly problem that is sufficiently representative of high-risk applications. For instance, stacking tasks are essential in biological gloveboxes, where stands and fixtures must be assembled for environmentally controlled experiments; and in manufacturing, where molds need to be assembled before metal casting. Second, the blocks in Jenga, with their elongated and thin shapes with wooden tones are similar to skin tones, making them interesting candidates to include in a challenging hand segmentation dataset. By demonstrating human-robot collaboration using such similar objects in our experiments, we increase confidence that our system can safely function in scenarios where humans and robots collaborate to handle more traditional SNM components.
@@
Green screen. A green screen was placed on the bottom of the glovebox for half of the participant videos. This green screen was used to later apply synthetic textures and colors to the background of the real-world images (see Fig. 2). The script incorporated in our code repository for generating synthetic images leverages DuckDuckGo Image searches. Notably, these images are not stored within our dataset; rather, we provide the capability for other researchers to independently construct analogous datasets. Frames extracted from videos containing a green screen background are placed into an OOD test set.
Gloves. Despite operating in a glovebox, half of the recorded videos feature ungloved hands (see the right side of Fig. 1.) We incorporated this additional OOD scenario due to the risk of glove tears. Thus, our dataset accounts for rare scenarios where the active safety system must still ensure safe HRC despite skin exposure. These videos are also placed in an OOD test set.
@@ 3.5 Data Annotation
Three classes were assigned to each image: left-hand, right-hand, and background. Human annotators were instructed to annotate each hand from the tip to the wrist, and provide their best estimate of the wrist location when the subject was wearing gloves. Using LabelStudio (Tkachenko et al., 2020), annotators provided key point prompts to MobileSAM (Zhang et al., 2023), a segmentation model which provided a coarse label for annotators to refine. Four of the researchers performed annotations. Two annotators labeled each image to develop inter-annotator agreement (IAA) for label quality. We calculate IAA in two ways: (1) as the average Cohen’s Kappa (0.916) and (2) as the average IOU (0.957) between annotator-provided labels across the full dataset , indicating strong agreement. Each frame’s annotations were converted to a single PNG file, where the three classes were recorded: left-hand, right-hand, background.
@@ 4 Experiments
We conducted three experiments to assess the challenging nature of our dataset. Experiment A is a transfer learning experiment designed to demonstrate the limitations of prior works, listed in Table 1, when applied to our task space. This experiment highlights the lack of sufficient transferability of pretraining on existing datasets to our specific domain. Experiment B studies uncertainty quantification (UQ) involving ID and OOD testing using the HAGS dataset. The OOD frames for all participants include scenarios where gloves are not worn or a green screen is placed in the background. For the green screen frames, we overlay images of different objects to create diverse and challenging testing conditions. This experiment assesses the robustness and reliability of our models in handling diverse, challenging scenarios. In Experiment C, we evaluate the skin tone diversity within industrial hand segmentation datasets using the Monk Skin Tone (MST) scale (Monk, 2019). The objective is to assess the effectiveness of these datasets in addressing the underrepresentation of diverse skin tones in machine learning research, thereby providing insights into their suitability for promoting inclusivity and fairness in model development.
@@ 4.1 Experimental Setup
For training all models in both experiments, we input 256x256 image sizes and use the Adam optimizer with a learning rate of 1e-3 for the UNet (30 million parameters) and BiSeNetv2 (5 million parameters) architectures, and 8e-4 for MobileSAM (10 million parameters) training. To train or perform inference with MobileSAM, a visual prompt is first required. We adopt a simple approach of selecting a bounding box prompt that contained the whole image. Dropout of p=0.1 was utilized for training of all architectures. We use a variety of data augmentations applied to our training set, including: resize, color jitter, advanced blur, Gaussian noise, and a random rotatation of 90 degrees (p=0.5).
Experiment A. To train, we use an internal cluster equipped with eight NVIDIA RTX A5000 GPUs. The models: UNet and BiSeNetv2 are trained as three member ensembles and undergo initial pretraining using the WorkingHands (WH) (Shilkrot et al., 2019), HaDr (Grushko et al., 2023), and HRC (Sajedi et al., 2022) datasets. Following this pretraining phase, we fine-tune the models on varying subsets of our ID HAGS dataset. This approach is designed to test the assertion of prior work that strong performance can be achieved with minimal real-world examples when leveraging their synthetic datasets. We perform the pretraining and fine tuning phases with early stopping on the validation step. We did not include the MECCANO (Ragusa et al., 2020) dataset in our study because it only provides bounding box labels, and our task requires pixel-level segmentation. We evaluate the models on both ID and OOD data from the HAGS dataset. Lastly, the transformer-based model MobileSAM was excluded from this experiment due to its high computational cost for training. This exclusion is not considered a significant limitation, as results from Experiment B demonstrated that its performance, measured by IoU, was substantially lower compared to UNet and BiSeNet when fine-tuned from scratch.
@@
Experiment C. This study examines skin tone diversity across existing datasets. The analysis involves the following preprocessing steps: (1) applying white balance, (2) normalizing illumination via histogram equalization, and (3) excluding shadow regions using a shadow threshold hyperparameter to the real world images in each dataset. After preprocessing, we isolate the region of interest corresponding to the ground truth mask for hands and skin. From this region, we compute the average RGB value of the pixels, convert it to a hexadecimal representation, and determine the closest match on the Monk Skin Tone Scale (Monk, 2019) (c.f. Figure 4(a)) using a minimum color distance metric. The matched tone is then assigned as the classification for that region. However, this method cannot be applied to the HaDr dataset because it does not provide ground truth masks for its real-world images, precluding its inclusion in the analysis. Lastly, for the HAGS dataset, we focus exclusively on the no glove, out-of-distribution (OOD) images because it is most pertinent to the objectives of this study.